# Causally constrained reduced-order neural models of complex turbulent dynamical systems

### v2

@@ Abstract
We introduce a flexible framework based on response theory and score matching to suppress spurious, noncausal dependencies in reduced-order neural emulators of turbulent systems, focusing on climate dynamics as a proof-of-concept. We showcase the approach using the stochastic Charney–DeVore model as a relevant prototype for low-frequency atmospheric variability. We show that the resulting causal constraints enhance neural emulators’ ability to respond to both weak and strong external forcings, despite being trained exclusively on unforced data. The approach is broadly applicable to modeling complex turbulent dynamical systems in reduced spaces and can be readily integrated into general neural network architectures.
@@ I Introduction
Modeling and understanding complex turbulent dynamical systems often requires abandoning full microscopic descriptions in favor of reduced-order models, where coarse-grained variables evolve according to effective stochastic dynamics [12, 67]. Such reduced stochastic descriptions are fundamental across physics, ranging from chemical physics [47, 88], neuroscience [41, 89], mathematical physics [91, 90], turbulence [66, 94], climate science [42, 64, 69, 34, 59, 50, 49], and many others. In practice, one is often faced with an inverse problem: given long trajectories of a Markovian, coarse-grained system, construct a data-driven model that faithfully reproduces its statistical and dynamical features [86, 43, 83, 73, 54, 51, 11, 98, 17, 16, 36, 3, 31, 29]. Recent advances in machine learning further contribute to inverse modeling by providing flexible neural network architectures for data-driven emulation, with promising applications in fluid dynamics [33, 75, 103], weather [81, 9], and climate science [105, 104, 40, 14, 24]. However, neural network–based models are typically agnostic to the underlying physics, which can lead to nonphysical behavior, such as violations of energy conservation - an issue that can be addressed through appropriate constraints [73, 51, 7, 77] - or incorrect responses to external forcings. The latter issue has been documented in neural weather models [101] and, especially, in climate emulators [102, 106]. This limitation highlights a central open challenge for neural emulators of physical systems: high skill in reproducing stationary statistics does not guarantee accurate responses to external perturbations, thereby limiting the use of such models for process understanding through sensitivity analysis and the exploration of “what if?” scenarios. Crucially, a model’s ability to respond to forcings is inherently linked to causality, since cause–effect relationships in physical systems are inferred through external interventions [46, 2, 60, 62], commonly realized via perturbation–response experiments (see [82, 8] in the contexts of neuroscience and climate, respectively). This connection suggests that constraining a data-driven model to respect a given “correct” causal structure should improve its ability to respond to external forcings. The resulting challenge is twofold: to infer causal relationships from stationary, unperturbed data alone, and to incorporate this information into neural emulators, all within a computationally efficient framework that is both scalable and practical. In this context, the Fluctuation–Dissipation Theorem (FDT) plays a fundamental role as it provides a bridge between forced and natural fluctuations of physical systems [52, 30, 63, 74, 39, 61]. The work of Baldovin et al. (2020) [5] showed that the FDT provides a rigorous causal inference strategy in the interventional sense directly from data, and this approach has since been adopted in a growing body of work [13, 4, 28, 27, 1, 36, 35, 29]. In this Letter, we take this line of research one step further by leveraging the FDT to constrain the training of general neural emulators of physical systems. Given a trajectory $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$ of a multivariate Markovian system, our framework proceeds in two steps: (i) use the FDT to infer the direct causal couplings $x_{j}\rightarrow x_{k}$ across degrees of freedom $x_{j}$, $x_{k}$; (ii) learn a stochastic neural model from $\mathbf{x}(t)$ while preserving the inferred causal structure. A related direction was recently proposed by Chen and Liu (2025) [16], who used causation entropy to rank candidate functions in the construction of causally constrained reduced-order models, highlighting the benefits of causally informed approaches. Our work addresses a complementary setting by providing a practical strategy to enforce a prescribed causal structure within general neural network architectures. To illustrate the methodology, we focus on climate dynamics as a representative example of a complex, turbulent dynamical system that is necessarily modeled through an effective, coarse-grained description [69, 51, 66], and for which accurate modeling of forced responses is essential [29]. We consider a stochastic version of the Charney–DeVore model [15], relevant for conceptual studies of low-frequency climate variability. This choice provides a simplified testbed that nevertheless retains key features of realistic coarse-grained climate dynamics, including strong mixing, fast decay of autocorrelations, and regular, unimodal distributions [70, 68]. We show that causally constrained neural emulators substantially improve the ability of neural networks to capture responses to both weak and strong forcings. In particular, the causal emulator accurately reproduces responses beyond the linear regime, including time-dependent shifts of the attractor in state space, despite being trained solely on stationary, unperturbed data.
II Framework: from linear response to causal constraints
Fluctuation–Dissipation Theorem (FDT) and causality
@@ II Framework: from linear response to causal constraints
Given an nn–dimensional stochastic nonlinear system $\mathbf{x}(t)$, the FDT [63, 74] states that the linear response, in the ensemble average sense, of a degree of freedom $x_{k}(t)$ to an infinitesimally small impulse perturbation $\delta x_{j}(0)$ applied to component $x_{j}(0)$ can be written as:
@@
$\begin{split}R_{k,j}(t)&=\frac{\delta\langle x_{k}(t)\rangle}{\delta x_{j}(0)}=-\langle x_{k}(t)s_{j}(\mathbf{x}(0))\rangle,\\ s(\mathbf{x})&=\nabla\log\rho(\mathbf{x}).\end{split}$ (1)
$\rho(\mathbf{x})$ is the system’s invariant density and $s(\mathbf{x})$ is the score function of the system [37, 35]. The focus on stochastic systems allows us to circumvent difficulties with determistic dissipative dynamical systems, where the invariant measure is singular [93, 92, 19]. The brackets $\langle\cdot\rangle$ represents ensemble averages, computed over the invariant density. Eq. (1) provides a rigorous strategy for inferring causal relations in physical systems in the interventional sense from data alone [5]. As discussed in [29], $R_{k,j}(t)$ supports multiple causal interpretations: values at the shortest time scale ($t=dt$) encode direct causal links $x_{j}\rightarrow x_{k}$. Nonzero responses at later times reflect indirect causal links. Additionally, a cumulative measure of causal influence can be defined as $D_{k,j}=\int_{0}^{\infty}R_{k,j}(t)dt$ [5]. The causal constraints proposed in this work focus on the direct causal links, $R_{k,j}(dt)$. In practice the response is computed for standardized data as $(\sigma_{j}/\sigma_{k})R_{k,j}(dt)$, $\sigma_{i}$ representing the standard deviation of $x_{i}$; for simplicity, we denote this standardized operator by $R_{k,j}$.
FDT from data
The main challenge in estimating responses via Eq. (1) lies in estimating the score function $s(\mathbf{x})$, particularly in high-dimensional settings. Traditional studies have adopted the quasi-Gaussian (qG, hereafter) ansatz [55, 63], in which $\rho$ is approximated by a Gaussian. The qG approximation has proven effective for predicting mean responses, even for nonlinear problems [63, 39, 70, 35]. Recent developments in score-based generative modeling [96, 95], offer new ways forward to overcome previous approximations by parametrizing the score function directly from data. Importantly, when provided with enough data, these new strategies scale to high-dimensional systems such as quasi-geostrophic flows [84], idelized ocean flows [97], weather [87], climate [10], fluid dynamics [56], among others. The estimated score $s(\mathbf{x})$ can then be inserted into the analytical response formula in Eq. (1), circumventing the need for explicit approximations of the invariant density, as in the quasi-Gaussian case. This approach was first proposed in [37]. These advances are leading to a set of score-estimation techniques rather than a single strategy. For example, [35] highlighted different denoising score-matching procedures tailored to the dimensionality of the underlying system. Here, we further contribute to this “FDT–generative modeling toolbox”. Specifically, we adopt the score-matching estimator proposed in [45], allowing us to infer the score directly on the attractor, rather than up to a small noise perturbation (e.g., as in denoising score matching). The estimated score is used as an initial approximation and is then refined using the correction procedure introduced in [37]. Details are reported in Section I of the Supplemental Material.
@@
Consider a stochastic dynamical system
$\dot{x_{i}}(t)=f_{i}(x_{1},x_{2},...,x_{n})+\sigma_{i}\xi_{i}(t),~\text{with}~i=1,2,...,n,$
where $\sigma_{i}\xi_{i}(t)$ is a stochastic forcing, $\bm{\xi}(t)$ being a standard Gaussian white noise process. A perturbation to $x_{j}$ can affect $x_{k}$ at the next time step only if $f_{k}$ depends on $x_{j}$. Thus, in the limit $dt\to 0$, $\partial f_{k}/\partial x_{j}\equiv 0$ implies the absence of a direct causal link and consequently, a vanishing short-time response $R_{k,j}(dt)=0$. In practice, the system’s Jacobian is not observable from data, and a reliable numerical estimation is unfeasible in high-dimensional systems. However, the short-time response $R_{k,j}(dt)$ can be efficiently estimated from unperturbed data using the FDT in Eq. (1) as shown in [35]. A vanishing response at the shortest resolved time scale thus implies $\big\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\big\rangle=0$ for $j\neq k$. This information on vanishing entries of the ensemble-averaged Jacobian is used to constrain neural architectures. Thus, given a long trajectory of $\mathbf{x}(t)$, we first estimate the response operator $\mathbf{R}(dt)\in\mathbb{R}^{n,n}$ with the score-matching procedure introduced earlier. We then proceed by constructing a binary adjacency matrix $\mathbf{A}\in\mathbb{R}^{n,n}$, by thresholding $\mathbf{R}(dt)$. $A_{k,j}=0$ denoting the absence of a direct causal link $x_{j}\rightarrow x_{k}$, and $A_{k,j}=1$ denoting its presence. One way to threshold $R_{k,j}(dt)$ is to examine its values on a logarithmic scale; a visual inspection of the response magnitudes often reveals a clear separation between significant links and noise. The absence of such separation may itself be informative, suggesting a densely connected system, in which case no additional causal constraints are needed. To provide an unsupervised criterion, we propose the following procedure:
@@
    Self-interactions. We assume that for any dissipative dynamical system $x_{j}\rightarrow x_{j}$ and therefore $A_{j,j}=1$.
@@
    Cross-interactions. We consider the off-diagonal values of $R_{k,j}(dt),j\neq k$, and apply a logarithmic transformation $\ln|R_{k,j}(dt)|$. This helps separate potential causal links (larger responses) from the bulk of spurious, near-zero interactions. A $k$-means algorithm [85] ($k=2$) is then used to partition these log-scaled values into two clusters, corresponding to “significant” and “insignificant” responses. An entry of $A_{k,j}=1$ is set if its corresponding $\ln|R_{k,j}(dt)|$ value belongs to the cluster with the larger centroid; otherwise, it is set to 0.
The causal Adjacency matrix is then used to constrain the training of a neural emulator. Given a $n$-dimensional trajectory $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$ we fit a neural operator $\mathbf{f}$ such that $\dot{\mathbf{x}}(t)=\mathbf{f}(\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$. The causal constraints are imposed on the deterministic drift $\mathbf{f}$. We learn $\mathbf{f}$ via a mean squared error (MSE) minimization such that $\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))$, where the time unit has been rescaled by the time sampling interval. Details on the learning process are found in the SM, Section III. A “vanilla”, unconstrained operator $\mathbf{f}$ is trained by minimizing MSE as $\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}))$. The causally constrained emulator is trained by minimizing the following loss function:
$\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}(t)))+\lambda\sum_{k,j\in\mathcal{S}}(\frac{\partial f_{k}}{\partial x_{j}})^{2}~,$ (2)
where $\mathcal{S}$ is the set of pairs $(k,j)$ for which the causal adjacency matrix $A_{k,j}=0$. $\lambda$ is a parameter determining the strength of the regularization. The derivatives are obtained via automatic differentiation [80]. We adopt soft constraints for two main reasons. First, they are computationally efficient [77]. Second, in real-world, high-dimensional climate systems with partial observations and finite data, causal estimates are inherently uncertain. In this regime, hard constraints based on imperfect causal information will introduce biases into the learned dynamics. Soft constraints, in contrast, provide a flexible form of regularization, ensuring an appropriate balance between data fidelity, as captured by the MSE, and causal regularization. Additionally, we note that the regularization in Eq. (10) can be adapted to novel architectures, including generative models [18], which we leave to future work.
@@
The proposed constraints leverage information about zeros of the ensemble-averaged Jacobian to constrain the state-dependent Jacobian. While this choice is practical and efficient, it can be limiting in idealized systems, for example because of exact symmetries. In such cases, nonzero local couplings may cancel in ensemble averages, yielding $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$ even if $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$, as occurs, for example, in the Lorenz ’63 model [57]. In such idealized settings, applying the proposed approach may require a symmetry-reduction step, as advocated in [22] and demonstrated in [76] for the Lorenz ’63 system. These limitations are not restrictive for realistic turbulent system considered here, such as weather and climate, which lack exact symmetries, are only partially observed, and admit multiple effective coarse-grained representations. There is no unique law to be discovered, instead the learned drift $\mathbf{f}$ represents one admissible effective realization, regularized to ensure physically consistent response behavior and agreement with observed statistics. Further details are provided in Section II of the SM.
Practical assessment of causal constraints from data alone.
Finally, the utility of causal constraints for a given system can be readily assessed by fitting a constrained model and an unconstrained counterpart. A key diagnostic is the comparison of the response operator computed from the emulators (see Section V of the SM) with the FDT-based operator in Eq. (1). When the inferred causal structure is informative, the constrained emulator improves the representation of linear responses without degrading the representation of stationary statistics. Moreover, comparing modeled and FDT-inferred response operators provides a natural criterion for a posteriori calibration of neural emulators based on forced responses. Related response-based calibration ideas have been explored in previous studies [71, 72]; a systematic formulation tailored to neural emulators and causal constraints is left for future work.
III Causal constraints in practice
Test-model for low-frequency climate variability.
To showcase the framework, we consider a stochastic version of the Charney–DeVore model [15], “CdV” hereafter. The model was obtained as a six-dimensional Galerkin truncation of the barotropic vorticity equation on a $\beta$-plane channel with topography. The formulation considered here is the one further proposed by De Swart [23]. The model has been widely studied in the context of large-scale atmospheric circulation [20, 21, 26], as well as in computational studies of regime transitions and rare events [38]. The model takes the following form:

### v3

# Physics and causally constrained discrete-time neural models of turbulent dynamical systems

@@ Abstract
We present a framework for constructing physics and causally constrained neural models of turbulent dynamical systems from data. We first formulate a finite-time flow map with strict energy-preserving nonlinearities for stable modeling of temporally discrete trajectories. We then impose causal constraints to suppress spurious interactions across degrees of freedom. The resulting neural models accurately capture stationary statistics and responses to both small and large external forcings. We demonstrate the framework on the stochastic Charney–DeVore equations and on a symmetry-broken Lorenz–96 system. The framework is broadly applicable to reduced-order modeling of turbulent dynamical systems from observational data.
@@ I Introduction
Modeling and understanding complex turbulent dynamical systems often requires abandoning full microscopic descriptions in favor of reduced-order models, where coarse-grained variables evolve according to effective stochastic dynamics [10, 59]. Such reduced stochastic descriptions are fundamental across physics, ranging from chemical physics [41, 79], neuroscience [35, 80], mathematical physics [82, 81], turbulence [58, 84, 14], climate science [36, 57, 61, 28, 52, 43, 42], and many others. In practice, one is often faced with an inverse problem: given long trajectories of a Markovian, coarse-grained system, construct a data-driven model that faithfully reproduces its statistical and dynamical features [77, 37, 74, 65, 46, 44, 9, 88, 15, 14, 30, 2, 25, 23]. Recent advances in machine learning further contribute to inverse modeling by providing flexible neural network architectures for data-driven emulation, with promising applications in fluid dynamics [27, 67, 94], weather [72, 6], and climate science [96, 95, 34, 12, 20]. However, neural network–based models are typically agnostic to the underlying physics. This can lead to several failure modes, including long-term nonphysical behavior, such as violations of energy conservation [65, 44, 69] and consequent blow-up of solutions [75], and incorrect responses to external forcings [90, 93, 97]. The latter limitation is particularly critical for scientific applications, where models are used for process understanding and exploration of “what if?” scenarios. Crucially, a model’s ability to respond to forcings is inherently linked to causality, since cause–effect relationships in physical systems are defined by their response to external interventions [40, 1, 53, 55], typically realized via perturbation–response experiments [73, 5]. In this Letter, we address these challenges by presenting an efficient strategy to construct physics and causally constrained neural models of turbulent dynamical systems from data alone. We focus on the less explored context of discrete-time modeling, which avoids inconsistencies arising in applications when continuous-time conservation principles are applied to temporally coarse-grained data. Given a trajectory $\mathbf{x}_{t}=(x^{(1)}_{t},x^{(2)}_{t},...,x^{(n)}_{t})$ of a multivariate discrete Markovian system, our framework proceeds in two steps: (i) Physics Constraints: We propose a neural model, with strict energy-conserving nonlinearities, tailored to realistic, discrete-time coarse-grained observations. (ii) Causal Constraints: We identify direct causal couplings $x^{(j)}\rightarrow x^{(k)}$ from data via the Fluctuation-Dissipation Theorem (FDT) and use this information to suppress spurious, non-causal interactions during model training.
The FDT plays a fundamental role here, as it provides a rigorous bridge between forced and natural fluctuations [45, 56, 54], allowing for a causal inference strategy in the interventional sense directly from unperturbed, stationary data [3, 11, 23]. Ultimately, the physics constraints, specifically formulated for discrete settings, provide the foundation of our approach. These physical priors promote robust inference in data-scarce regimes and ensure long-term numerical stability. Furthermore, this structural rigidity naturally bounds the impact of imperfect causal estimates: even if the inferred causal links contain errors, the resulting model remains confined to a physically consistent solution. As a representative application, we consider climate dynamics, where accurate prediction of forced responses is essential [23]. We test the framework on two benchmarks: a stochastic version of the Charney–DeVore model [13] and a symmetry-broken Lorenz–96 system [51] as a higher-dimensional testbed. The resulting neural emulators are stable, capture stationary statistics, and reproduce responses to both weak and strong external forcings, including time-dependent shifts of the attractor, despite being trained exclusively on unperturbed data.
II Framework: physics and causal constraints
II.1 Finite-time flow map with energy-preserving nonlinearities
Previous work.
We build on the abstract representation of turbulent dynamical systems often adopted by Majda and collaborators [56, 59] $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{A}\mathbf{x}+\mathbf{B}(\mathbf{x},\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, which arises when a large class of fluid flows are projected onto orthogonal basis functions. The quadratic term is energy-conserving, satisfying $\mathbf{x}\cdot\mathbf{B}(\mathbf{x},\mathbf{x})=0$ [50, 47], with the system’s energy defined as $E=(1/2)\mathbf{x}^{\mathrm{T}}\mathbf{x}$. $\mathbf{F}$ represents external forcing, and the linear operator $\mathbf{A}$ is typically decomposed as $\mathbf{A}=\mathbf{L}+\mathbf{D}$, where $\mathbf{L}$ (skew-symmetric) and $\mathbf{D}$ (symmetric, negative-definite) represent dispersion and dissipation, respectively. $\mathbf{\Sigma}\bm{\xi}(t)$ is a stochastic forcing, with $\bm{\xi}(t)$ denoting a standard Gaussian white-noise process. Following recent work on neural closure and modeling [92, 84, 83, 30], we replace $\mathbf{B}(\mathbf{x},\mathbf{x})$ by a general nonlinear operator $\mathbf{S}(\mathbf{x})\mathbf{x}$, where $\mathbf{S}(\mathbf{x})$ is constrained to be skew-symmetric, $\mathbf{S}^{\mathrm{T}}(\mathbf{x})=-\mathbf{S}(\mathbf{x})$. This preserves the key physical property $\mathbf{x}^{\mathrm{T}}\mathbf{S}(\mathbf{x})\mathbf{x}=0$, so the nonlinear term redistributes energy across modes without net injection or dissipation. The importance of energy-conserving nonlinearities in data-driven models of turbulent systems was highlighted in [65]. Details are presented in the Supplemental Material (SM) Sec. I.
Finite-time flow map formulation.
Observational data are fundamentally discrete and temporally coarse-grained, yet physics constraints are typically defined strictly in the continuous limit. To resolve this discrepancy, we introduce a discrete, abstract formulation of turbulent dynamical systems with exact energy-conserving nonlinearities for arbitrary time increments:
$\mathbf{x}_{t+1}=\mathbf{f}(\mathbf{x}_{t})=\mathbf{F}+\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}+\mathbf{\Sigma}\bm{\xi}_{t},$ (1)
where $\mathbf{F}$ and $\mathbf{\Sigma}\bm{\xi}_{t}$ represent the effective deterministic and stochastic forcings integrated over the finite time step. The nonlinear operator $\mathbf{Q}(\mathbf{x}_{t})$ is constrained to be strictly orthogonal ($\mathbf{Q}^{\top}\mathbf{Q}=\mathbf{I}$). Thus, the deterministic map splits into an energy-preserving rotation, $\mathbf{v}_{t}=\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$, followed by a linear transformation $\mathbf{M}\mathbf{v}_{t}$. The isolated nonlinear term contributes no net growth or decay to the energy budget, strictly preserving the $L_{2}$-norm of the state:
$\|\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}\|^{2}=\mathbf{x}_{t}^{\top}\mathbf{Q}^{\top}(\mathbf{x}_{t})\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}=\|\mathbf{x}_{t}\|^{2}.$
Consequently, energy growth in the forward map is bounded by the dominant singular value of $\mathbf{M}$. While additional constraints on $\mathbf{M}$ could be imposed, they are often not required in practice and may degrade performance. In the SM (Sec. II), we provide mathematical justification for Eq. (1). First, we demonstrate that in the continuous limit this mapping arises as a Lie–Trotter splitting scheme [7], and we further derive the corresponding effective deterministic and stochastic forcings $\mathbf{F}$ and $\mathbf{\Sigma}\bm{\xi}_{t}$. As illustrative examples, we recast the deterministic Lorenz-63 system and a stochastic triad model within this framework. Second, we return to the data-driven setting and show that the proposed discrete formulation in Eq. (1) enables stable learning from severely sub-sampled data, where both numerical integrators and continuous-time data-driven models fail.
Neural model fitting.
Given a stationary trajectory $\mathbf{x}_{t}$, we parametrize the deterministic drift $\mathbf{f}(\mathbf{x}_{t})$ by training a multilayer perceptron (MLP) to minimize the mean squared error $\text{MSE}(\mathbf{x}_{t+1},\mathbf{f}(\mathbf{x}_{t}))$. States are standardized to zero mean and unit variance. We enforce the orthogonality of $\mathbf{Q}(\mathbf{x}_{t})$ by predicting the matrix exponential $\mathbf{Q}(\mathbf{x}_{t})=\exp(\mathbf{S}(\mathbf{x}_{t}))$ of a skew-symmetric matrix $\mathbf{S}(\mathbf{x}_{t})$. Crucially, the linear terms are initialized with a “first guess” ordinary least squares (OLS) solution, while the MLP is initialized to yield $\mathbf{Q}(\mathbf{x}_{t})=\mathbf{I}$. This ensures the joint optimization starts from a Linear Inverse Model (LIM) [77] baseline before learning the nonlinear corrections. Here, the diffusion matrix $\mathbf{\Sigma}$ is estimated as a diagonal matrix from the standard deviation of the determistic residuals; the general framework captures spatial correlations via the Cholesky decomposition of the residual covariance. Further implementation details, including careful data rescaling, are discussed in SM Sec. I.
II.2 Fluctuation–Dissipation Theorem (FDT) and causality
@@
Given an nn–dimensional stochastic nonlinear system $\mathbf{x}_{t}=(x^{(1)}_{t},x^{(2)}_{t},...,x^{(n)}_{t})$, the FDT [56, 66] states that the linear response, in the ensemble average sense, of a degree of freedom $x^{(k)}_{t}$ to an infinitesimally small impulse perturbation $\delta x^{(j)}_{0}$ applied to component $x^{(j)}_{0}$ can be written as:
@@
$\begin{split}R^{k,j}_{t}&=\frac{\delta\langle x^{(k)}_{t}\rangle}{\delta x^{(j)}_{0}}=-\langle x^{(k)}_{t}s_{j}(\mathbf{x}_{0})\rangle,\\ s(\mathbf{x})&=\nabla\log\rho(\mathbf{x}).\end{split}$ (2)
$\rho(\mathbf{x})$ is the system’s invariant density and $s(\mathbf{x})$ is the score function of the system [31, 29]. The brackets $\langle\cdot\rangle$ represents ensemble averages, computed over the invariant density. Eq. (2) provides a rigorous strategy for inferring causal relations in physical systems in the interventional sense from data alone [3, 23]. Values at the shortest time scale $R^{k,j}_{1}$ encode direct causal links $x^{(j)}\rightarrow x^{(k)}$ and are used to construct our causal constraints. Evaluating Eq. (2) requires the score function $s(\mathbf{x})$. We leverage recent advances in score-based generative modeling [86, 85, 75, 87, 78, 8, 49] to learn $s(\mathbf{x})$ directly from data for use in the analytical FDT formula [31, 29], therefore bypassing traditional quasi-Gaussian approximations [48, 56]. Specifically, we employ the exact score-matching estimator [39] to evaluate the score directly on the system’s attractor. The resulting estimate is further refined using the correction procedure of [31] (see SM Sec. III for details).
@@
Given a discrete stochastic systems $x^{(i)}_{t+1}=f_{i}(x^{(1)}_{t},x^{(2)}_{t},...,x^{(n)}_{t})+\sigma_{i}\xi^{(i)}_{t}~\text{with}~i=1,2,...,n$, a perturbation to $x^{(j)}$ can affect $x^{(k)}$ at the next time step only if $f_{k}$ depends on $x^{(j)}$. Thus, $\partial f_{k}/\partial x^{(j)}\equiv 0$ implies the absence of a direct causal link and consequently, a vanishing short-time response $R^{k,j}_{1}=0$. In practice, the system’s Jacobian is not observable from data, and a reliable numerical estimation is unfeasible in high-dimensional systems. However, the short-time response $R^{k,j}_{1}$ can be efficiently estimated from unperturbed data using the FDT in Eq. (2). A vanishing response at the shortest resolved time scale thus implies $\big\langle\partial f_{k}(\mathbf{x})/\partial x^{(j)}\big\rangle=0$ for $j\neq k$. This information is used to constrain neural architectures. Thus, given a long trajectory of $\mathbf{x}(t)$, we first estimate the short-time response operator $\mathbf{R}_{t=1}\in\mathbb{R}^{n,n}$. We then construct a binary adjacency matrix $\mathbf{A}\in\mathbb{R}^{n,n}$, by thresholding $\mathbf{R}_{1}$. $A^{k,j}=0$ denoting the absence of a direct causal link $x^{(j)}\rightarrow x^{(k)}$, and $A^{k,j}=1$ denoting its presence. A visual inspection of $R^{k,j}_{1}$ on a logarithmic scale often reveals a clear separation between significant links and noise. The absence of such separation suggests a densely connected system with no need for causal constraints. To provide an unsupervised criterion, we propose the following procedure:
@@
    Self-interactions. We assume that for any dissipative dynamical system $x^{(j)}\rightarrow x^{(k)}$ and therefore $A^{j,j}=1$.
@@
    Cross-interactions. We consider the off-diagonal values of $R^{k,j}_{1},j\neq k$, and apply a logarithmic transformation $\ln|R^{k,j}_{1}|$, separating potential causal links (larger responses) from the bulk of spurious, near-zero interactions. A k-means algorithm [76] (k=2) is used to partition these values into two clusters. An entry of $A^{k,j}=1$ is set if its corresponding $\ln|R^{k,j}_{1}|$ value belongs to the cluster with the larger centroid; otherwise, it is set to 0 (absence of causal link $x^{(j)}\rightarrow x^{(k)}$).
The matrix $A^{k,j}$ is then used to constrain neural emulators. Specifically, the MSE minimization is regularized with a causal penalty:
$\mathcal{L}$ $=\mathcal{L}_{\text{MSE}}+\mathcal{L}_{Causal}$ (3)
$=\text{MSE}(\mathbf{x}_{t+1},\mathbf{f}(\mathbf{x}_{t}))+\lambda\sum_{k,j\in\mathcal{S}}(\frac{\partial f_{k}}{\partial x^{(j)}})^{2}~,$
where $\mathcal{S}$ is the set of pairs $(k,j)$ for which $A_{k,j}=0$. $\lambda$ is a parameter determining the strength of the regularization.
Imperfect causal estimates.
False negatives in the adjacency matrix $A^{k,j}$ (true links penalized as noncausal) can degrade the learned emulator when learning via Eq. (3). Physics-constrained architectures mitigates the impact of imperfect causal estimates through rigid structural priors. Additionally, if false negatives are suspected, via validation with stationary and perturbed statistics (see paragraph below), a more robust penalty can be employed:
$\mathcal{L}_{Causal}^{Robust}=\lambda\sum_{k,j\in\mathcal{S}}\min\!\Big((\frac{\partial f_{k}}{\partial x^{(j)}})^{2},\gamma\Big),$ (4)
which bounds the penalty once $(\partial f_{k}/\partial x^{(j)})^{2}\geq\gamma$, allowing the MSE term to dominate when interactions are mistakenly excluded. An heuristic estimate for $\gamma$ based on the FDT operator $R^{k,j}_{1}$ is given in Sec. IC in the SM.
@@
Causal constraints rely on zeros of the ensemble-averaged Jacobian to restrict the state-dependent Jacobian. In idealized systems with exact symmetries (e.g., the Lorenz ’63 system [50]), nonzero local couplings may cancel in the average, yielding $\langle\partial f_{k}(\mathbf{x})/\partial x^{(j)}\rangle=0$ even when $\partial f_{k}/\partial x^{(j)}\neq 0$. Such cases require a symmetry-reduction step [18, 68]. However, this is rarely restrictive for realistic, partially observed turbulent flows (e.g., weather and climate), which lack exact symmetries and admit multiple coarse-grained representations, see Sec. IV in the SM for details.
Assessing causal constraints.
Unlike rigid physical priors, the utility of causal regularization is system-dependent and contingent on the accuracy of the data-driven causal discovery. The benefit of augmenting physics-constrained models with causal information can be assessed directly from data by comparing the FDT-based operator (Eq. (2)) with the emulators’ response operators (SM Sec. V). When the inferred causal structure is informative, the resulting emulator improves linear response predictions without degrading stationary statistics.
III Physics and causal constraints in practice
We first illustrate the proposed framework on a six-dimensional stochastic system, highlighting the causal inference procedure and the benefits of the imposed constraints. We then focus on a symmetry-broken Lorenz–96 system as a more demanding test -case, evaluating both its response to strong external forcings and robustness against imperfect causal estimates.
III.1 Stochastic Charney-DeVore model
We consider a stochastic version of the Charney–DeVore model [13], “CdV” hereafter, later reformulated by De Swart [19]. The model was obtained as a six-dimensional Galerkin truncation of the barotropic vorticity equation on a $\beta$-plane channel with topography. The model has been studied in the context of large-scale atmospheric circulation [16, 17, 21, 32] and takes the following form: